OpenStudy (anonymous):

together lauren and aaron sold 456 raffle tickets. lauren sold 24 more than aaron. how many did lauren sell?

OpenStudy (anonymous):

456/2 = 228 That is the number if they both sold the same amount of tickets. Lauren sold 228. Aaron sold 228. However, Lauren sold 24 more so: Lauren sold 228+14 = 242 Aaron sold 228-14 = 214